A body starts from rest and moves in a straight path with inform acceleration of 8m/s2 for 5meters , it then decelerate uniformly to rest in the next 10 seconds



1
Expert's answer
2021-01-27T09:30:52-0500

 So, it is necessary to find the acceleration with which the body slowed down.


"s=\frac{v_f^2-v_0^2}{2a}\to v_f=\sqrt{2as}=\sqrt{2\cdot8\cdot5}=8.94\ (m/s)"


"v=v_f-a_1t\to \ a_1=v_f/t=8.94/10=0.894\ (m/s^2)" . Answer